OpenStudy (anonymous):

use a system of equations in two variables solve the following problem: 3 times a first number increased by twice a second number is 21. The difference between the first number and the second number is 2. Find the number?

OpenStudy (anonymous):

3x+2y=21 x-y=2 would be the equations, with x as the first number and y the second number. But 'Find the number' does not tell us which number you desire.

OpenStudy (anonymous):

If you multiply the second equation by 2 then 3x+2y=21 2x-2y=4 and add 5x=25 therefore x=5, subing into the second equation y=3 So the first number is 5 and the second number is 3.

OpenStudy (anonymous):

thanks for the steb by step